so im working on my homework and the equation is slope intercept and it's so confusing. i tried doing it okay, m=1/4 (-2,-3) than i got y-2=1/4(x--3) so you add two to three than 1/4 times x= 1/4x=14? than i can't figure out how to turn that into slope-intercept form.
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

so im working on my homework and the equation is slope intercept and it's so confusing. i tried doing it okay, m=1/4 (-2,-3) than i got y-2=1/4(x--3) so you add two to three than 1/4 times x= 1/4x=14? than i can't figure out how to turn that into slope-intercept form.
-----
You made a few mistakes.
m = 1/4
(-2,-3)
Form: y-k = m(x-x)
y + 3 = (1/4)(x+2)
---
y + 3 = (1/4)x + (2/4)
---
y+3 = (1/4)x + 1/2
---
y = (1/4)x +(1/2)-3
---
y = (1/4)x -(5/2)
---
slope = 1/4
intercept = -5/2
-------
An easier way to work the problem:
m = 1/4; (-2,-3)
Form y = mx+b
Substitute for m, x and y to solve for "b":
-3 = (1/4)(-2) = b
-3 = -1/2 + b
b = -5/2
----
Equation:
y = (1/4)x-(5/2)
=====================
Cheers,
Stan H.
===========

Answer by mananth(16949)

m=1/4 (-2,-3)
...
you have to write the equation in the form y= mx+b
First you have to find the value of b
..
plug value of m and the point(-2,-3)in the equation
..
-3=(1/4 *-2)+b
-3=-1/2 +b
add 1/2
-3+1/2 =b
-5/2 = b
..
Now form the equation
y = 1/4*x -5/2
OR in decimal form
y = 0.25x-2.5
..
m.ananth@hotmail.ca